History of the Islamic Middle East III The Armenian Massacres from 1894-1918

Dustin Mitchell March 26, 2001

1 Introduction

The history of Armenia and her people stretch back to the depths of history, over 3,000 years ago. Located between the Black Sea and the Caspian Sea, the nation possesses desirable natural resources. It also straddles an important land-bridge between Anatolia and the Northern territories that would eventually become Russia. During the period we are considering, Armenia did not have an independent existence. The Armenian people were divided between the Russian Empire and the Ottoman Empire. The Russian Armenians were treated reasonably well, but the Ottomans were extremely cruel in their treatment of their Armenian population.

Neither empire was interested in ruling the Armenian people themselves; both had designs on the land for other purposes. The Russians did not want a non-Russian presence in such a crucial part of the Empire. The Ottomans had no use for a Catholic people in an empire representing the Islamic world. Thus, from the mid-19th Century onward, it became increasingly clear to the international community that both empires wanted “Armenia without the Armenians”, and were prepared to use the cruelest of methods to obtain their goal. As these methods were put into practice, the Great Powers were slow to react, largely because of the restrictive environment of alliances and treaties binding them to one another.

2 1850-1896

Throughout the 19th Century, the Ottoman Armenians had been treated badly by the Empire. In 1969, they were subject to corrupt local Ottoman administration, unfair taxation, denial of justice, and a practice, called Kishlak, whereby the Armenians were forced to quarter the Kurds during the winter. Thus when the Russian Empire took large portions of Ottoman Armenia in the Russo-Turkish war (1877-78), securing proper treatment of the Armenians provided a convenient justification for continued occupation of the territories. After ceasing hostilities, Turkey and Russia signed the Treaty of San Stefano, which stated that the Russian troops would withdraw only after the Ottomans had implemented “the improvements and reforms demanded by local requirements in the provinces inhabited by the Armenians, and to guarantee their security from the Kurds and the Circassians” [?, 7].

Understanding that the Russians intended to retain Armenia indefinitely, the Ottoman Sultan appealed to the British, eventually surrendering Cyprus in exchange for protection from the Great Power. Only a few months after the Treaty of San Stefano, the Great Powers signed the Treaty of Berlin, under which the Russians would withdraw immediately, leaving the implementation of reforms in Turkish Armenia to the joint supervision of the Great Powers. For reasons discussed below, this supervision was ineffective in bringing about any reforms. Indeed, treatment of the Armenians continued to deteriorate.

The Sultan created a new force of Kurds with explicit instructions to create disturbances and depopulate the Armenian territories, settling Muslim Kurds in their place. By the 1880’s, the imperial administrators were no longer making any pretense of preventing the depredations of the Kurds, who were robbing, attacking, and killing Armenians in massive proportions. Taxation, as well as unofficial protection payments made to the Kurds, had driven many Armenians to leave for other lands, which was precisely the desired result.

In June of 1894, the peasants of Talori offered to pay the official taxes in full if the exactions of the Kurds were stopped. The local governor answered their proposal with violence, which quickly escalated. Soon the military intervened, resulting in massive bloodshed. Although there was some Armenian responsibility for the violence, firsthand accounts have since made it clear that the Ottoman reaction was far out of proportion to the original crimes. The Ottomans took great lengths to cover their misdeeds, claiming that only several hundred active and armed rebels had been killed, when indeed one of the commanding officers communicated to an American consul that the deaths numbered about 5,000, and were not limited to armed rebels.

Conflict continued, with several other similar incidents occurring in Turkish Armenia. On 26 August, 1896, a group of Armenians captured a European bank in Constantinople with the aim of drawing the attention of the great powers to their cause. Negotiations were held, and the rebels were given free passage out of Turkey. Protests followed throughout the capital, covertly instigated by the Ottoman authorities, and for two days Armenians throughout Constantinople were beaten mercilessly both by police and by ruffians given encouragement by the government.

The Treaty of San Stefano had given the Armenian people hope for better treatment. Russia’s motives, however, had not been humanitarian; with control of Turkish Armenia, Russia had access to warm-water ports, and found itself significantly closer to taking Constantinople, which would give the empire great power in the Eastern Mediterranean. It was precisely for this reason, however, that Britain was opposed to the conditions of the treaty. The British government sided with the Ottoman Empire, convening the Cyprus Convention which soon led to the creation of a new peace treaty in Berlin. It is significant to note that the Armenians were not invited to these conventions, and that the British interests were focussed on removing power from the Russian empire, not on assisting the Armenians.

Supervision by the Great Powers essentially consisted of “reminders” sent to the Sultan that he had promised meaningful reforms in Armenia. With the alliance system in place among the Powers, it was out of the question that any military action be taken to reinforce these reminders. Any action would require the agreement of all of the powers, and Russia and France were resistant to any attempt to interfere in the Ottoman Empire. Britain, although interested in more definitive action, had a great financial investment in the fragile Ottoman Empire, and was hesitant to take any steps that might lead to its demise.

Witnessing the continuation of the atrocities against the Armenians, the British Foreign Minister, Lord Salisbury, made an effort to demonstrate the naval power of the European nations. He was successful in securing the support of Germany, Austria-Hungary, and Italy, but he could not sway Russia and France from their hands-off policy. Russia, however, continued to claim reliance on the Sultan to introduce reform. The Russian Tsar expected that he would eventually possess Armenia, and if the Turks had cleared the region of Armenians then the Russians could easily settle it with their own peoples; he would have Armenia without the Armenians [?, 23].

By 1896, it was clear to the Sultan that the Great Powers would not and could not interfere with his annihilation of the Armenians. Although it was to be 13 years until a single incident would again attract such attention, the day-to-day attacks on the Armenians only increased.

3 1896-1914

Abdul Hamid, the ‘red Sultan’, was certainly aware of, if not responsible for, the atrocities during these painful years. Thus the Armenians were thrilled at the rise to power of the Young Turks in 1908. This new group, embodied in power by the CUP, espoused ideals of reform and equality, and set about making generous improvements the empire; Hamidian censorship was removed, education was expanded, government was streamlined, and the military was reformed. The Armenians were relieved of direct governmental persecution, and had hopes that law and order would be enforced, restricting the predations of the Kurds. All of this was done in the interest of Ottomanism — the equality of all Ottoman subjects regardless of nationality or religion. However, it quickly became apparent that Ottomanism would not be effective:

According to the constitution all subject people, Muslims and Christians, are equal before the law. But you one and all know that this is an unrealisable ideal … There can therefore be no question of equality, until we have succeeded in our task of Ottomanising the Empire. [?, 39]

In this context, Ottomanization was a policy of ruthless Turkification and Islamization of the subject peoples. Although minority religions were not prohibited from worshiping, they were forced to use the Turkish language and denied the right to organize. Returning to the laws of the Sharî’ah, Christians were again placed in a position legally subordinate to that of Muslims.

In 1909, the Turkish government took action to suppress what it claimed was a revolt in Cilicia begun by an Armenian political party, the Hunchakian party. However, a British observer claims that the official response was not localized to the place of the original disturbance, and that it began at about the same time everywhere. This suggests that it was not a response to a single uprising, but a premeditated attack on the Armenian population, very much like those under Hamid. Estimates of the death toll range up to 30,000, including not only men but women and young children.

Curiously, Sultan Abdul Hamid happened to be in power for the several days of this incident. Although the CUP was quick to blame the violence on the Sultan, it is more likely that the CUP instigated the violence to secure for themselves a quick return to power [?, 42-43].

Throughout this period, the Great Powers discussed the issue of the Armenians at length, but in a broader sense than before. Instead of searching for a way to alleviate the suffering of the Armenians within Turkey, the Powers discussed alternate solutions. The solution advocated by the Armenians was a new, sovereign state. Several years earlier, an optimistic Armenian had said, “Armenians are determined to be free. Europe listened to the Bulgarian horrors and made Bulgaria free. She will listen to our cry … We are desperate, we shall do it” [?, 16].

Russia was certainly not interested in such a development. A new state between the Russian Empire and the Bosphorus would further impede the empire’s efforts for Mediterranean access. The Tsar also worried that an independent state to replace Turkish Armenia might encourage nationalistic sentiments in Russian Armenia, which might then join the southern Armenians in their new nation. Russia instead advocated reform within the Turkish state; this was a complete reversal of their earlier opinion. This tactic had several advantages. As before, Russia desired access to the warm waters of the Black Sea, and with the popular support of the Turkish Armenians military passage through the region would likely be fairly easy.

Russia considered gaining this popularity by covertly stirring the Turkish Kurds into action, then occupying the region to “rescue” the Armenians. However, Russian intentions were all too plain to Turkey and the other Powers, who would not have allowed such a plan to succeed.

The British had come to the conclusion by this time that a Russian occupation of the Armenian territories would be best for the Armenians, but could not permit such a move. As before, Britain had strong economic reasons to maintain the unity of Turkey, and to maintain control of the Mediterranean.

While European and Russian leaders were debating a solution to the situation of the Armenians, the wholesale extermination of the Armenians in Turkey continued, with the encouragement of the government in Constantinople. Finally, in 1914, the Great Powers forced the Young Turks to accept an agreement whereby Turkish Armenia was divided into two administrative regions, each controlled indirectly by the Powers. It seemed that this agreement would finally put to rest the issue of Armenian livelihood in Armenia, but the Great War broke out before it could be put into action.

4 1915-1918

Only days before the outbreak of the Great War, Turkey chose to side with the Central Powers. The country was soon called into action by Berlin, and delivered a crushing blow to the Russians by closing the Straits to Allied traffic, effectively isolating Russia. Armenians on both sides of the border loyally fought alongside the soldiers of their respective empires. Despite this, in 1915, the Allies were defeated in their attempt to force the Dardanelles, the Turks took the resulting respite to eliminate their Armenian problem permanently.

Although the Armenians were extremely loyal throughout the history of the empire, particularly in defending the empire from the Allied powers in the Great War, the CUP claimed that there had been a revolt of disloyal Armenians at Van, on the Turkish side of the border with Russia. As a security measure, the government ordered the immediate deportation of all Armenians in the region. Even allowing that a disturbance may have occurred at Van, deportations were begun several days before the date given for the revolt. Leslie Davis, an American Consul in Harput describes the deportation there:

There were perhaps 3,000 who left Mamouret-ul-Aziz on that hot July day. A few had ox-carts, a few had mules or donkeys, some carried their scanty effects on cows, but the most of the people had to leave on foot carrying their baggage on their backs and their children in their arms on the terrible journey over the burning sands of the desert. The return soon afterwards of the donkeys and mules which the Government had provided indicated all

too well the fate of those who had left.[?, 59]

The Armenians were marched great distances over desert lands in the heat of Summer, given little or no water or food, preyed upon by criminals released from jail for this purpose. In Constantinople, Armenians of all ages were executed, sold as slaves, or sent on these terrible death-marches.

Descriptions of the horrors are plentiful, and painful to read. The empire had recently declared itself free of the Capitulations, which had offered some protection for Christians, and the CUP had just declared a jihad, so national sentiments were certainly not in favor of the Armenians. It seems that the Turks were using the cover of the brutalities of war to finally achieve their cruel goal of the elimination of all Armenians in Armenia.

Despite an official declaration from Turkey that the movements had been a routine matter of internal security, the Russian, British, and French governments were not long in reaching the conclusion that the Turks had acted with extreme cruelty. However, being already at war with Turkey, they were powerless to act. Similarly, Germany and Austria were could not destroy their alliance with Turkey for the sake of the Armenians.

One of the three leaders of the CUP, Djemal Pasha, expressed willingness to take over the administration of Turkey, and entered involved discussions with the Allied Powers on this topic at the end of 1915. He offered to sacrifice the Dardanelles and Constantinople and to properly treat the Armenians, in exchange for active support from the Allies against the existing Turkish government. The European powers, except France, were eager to implement this plan, but because of French resistance it never came to fruition.

Finally, the Sykes-Picot agreement and the post-War treaties broke the old Ottoman Empire up, dividing the Armenian territories between France and Russia and ending Turkish persecution of the Armenians.

5 Conclusion

Despite real concern for the welfare of the Armenian people in Turkey, the European nations were unable to bring about any significant improvement in their living conditions. Although the issue was a significant one in world politics at several points during the period of concern, at all times the needs of the Powers outweighed the needs of the Armenian people.

It is difficult to lay blame on any of the European nations in this episode. Although the conditions of the Armenians were terrible, in many cases the Powers made legitimate decisions in the balancing of their needs. For instance, if Russia had proceeded to annex Turkish Armenia in 1909, the system of alliances throughout Europe would likely have begun the Great War four years early.

On the other hand, there are many situations parallel to this one in the present age. The world community is aware of gross abuses of human rights in many areas, but if and when some nation intercedes, the actions taken are colored by the national requirements of that nation, and are often not the best solution to the problem.